OpenStudy (ashking1):

Solve the proportion x/x+1=4/x+4 . What is the value(s) of x

Parth (parthkohli):

First you cross multiply. \( \color{Black}{\Rightarrow x(x + 4) = 4(x + 1) }\)

jhonyy9 (jhonyy9):

why x^2 +1=0 ? i think x^2 -4=0

jhonyy9 (jhonyy9):

so than x^2 =4 x_1,2=+/- 2

OpenStudy (ashking1):

the choices were x=-2 x=4 x=-2 and x=2 x=-1 and -4

Parth (parthkohli):

Solve using the product rule

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - 4 = 0 }\) \( \color{Black}{\Rightarrow (x + 2)(x - 2) = 0 }\)

Parth (parthkohli):

lol I was wrong twice ;)

jhonyy9 (jhonyy9):

on the right part not is 4x+1

Parth (parthkohli):

Yeah I got that wrong but deleted it

jhonyy9 (jhonyy9):

so than x_1,2=+/- 2

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + 2 = 0 }\) \( \color{Black}{\Rightarrow x = -2 }\) OR \( \color{Black}{\Rightarrow x - 2 = 0 }\) \( \color{Black}{\Rightarrow x = 2 }\)

OpenStudy (ashking1):

ok thks

jhonyy9 (jhonyy9):

yw good luck bye